# Anomalous current fluctuations from Euler hydrodynamics

### v1

@@ Abstract
We consider the hydrodynamic origin of anomalous current fluctuations in a family of stochastic charged cellular automata. Using ballistic macroscopic fluctuation theory, we study both typical and large fluctuations of the charge current and reproduce microscopic results which are available for the deterministic single-file limit of the models. Our results indicate that in general initial fluctuations propagated by Euler equations fully characterize both scales of anomalous fluctuations. For stochastic dynamics, we find an additional contribution to typical fluctuations and conjecture the functional form of the probability typical distribution, which we confirm by numerical simulations.
Introduction.—Statistical physics aims to describe universal behaviors emerging from interactions of a large number of microscopic degrees of freedom. While traditionally the main focus has been equilibrium physics, in recent years it has been appreciated that out-of-equilibrium properties of the system can also display rich universal behaviors. A prominent tool that describes such physics is macroscopic fluctuations theory (MFT) [1], which captures large fluctuations of driven diffusive systems. The validity of MFT has been verified for a variety of systems [2, 3, 4, 5], but large fluctuations in these systems usually behave in a regular way with finite scaled cumulants of observables. This naturally raises the question of whether there exists systems that break such regularity and support new dynamical phenomena.
Recent numerical studies have found dynamical criticality and anomalous spin fluctuations in integrable spin chains [6, 7], which have also been probed experimentally using superconducting qubits [8]. While the mechanism responsible for anomalous spin fluctuations remains to be clarified, an exact calculation of fluctuations in a deterministic charged cellular automaton [9] provided a basic example of dynamical criticality in a deterministic many-body system. The procedure has been generalized to a wider class of single-file dynamics, which was explored in Refs. [10, 11, 12, 13, 14, 15]. Furthermore, anomalous fluctuations have also been observed in the easy-axis regime of an integrable spin chain [7, 16] and in Dirac fluids [17], hinting at a more general setting supporting anomalous fluctuations.
@@
A system of three-mode hydrodynamic equations similar to Eq. (3) along with the free chiral equations for ϱ±subscriptitalic-$\varrho_{\pm}$ was also recently derived for Dirac fluids by keeping only the most relevant terms based on symmetry considerations in Ref. [17]. In the present case, the hydrodynamic equations are exact (i.e. no terms that contribute at the Euler scale are omitted), and interestingly, are completely insensitive to the crossing probability $\Gamma$. That said, as we will see later, typical fluctuations do turn out to depend nontrivially on $\Gamma$. Before discussing this dependence, we first study large fluctuations of the SCCA, which can be fully characterized using only Euler hydrodynamic data.
@@
As was also observed in Ref. [9], the SCGF Eq. (15) is anomalous in that the scaled higher cumulants $c_{2n>4}=\lim_{t\to\infty}t^{-1}\langle[\hat{J}(t)]^{2n}\rangle^{c}$ are singular. BMFT clarifies the physical origin of this anomalous behavior: since the normal mode $b(x,t)$ of the charge $\varrho_{c}(x,t)$ travels along the ray $x=0$, fluctuations along $x=0$ proliferate and thereby render clustering of connected charge current correlations algebraic in time [25]. While such a scenario generally results in superdiffusion [26], it is avoided here due to the charge conjugation symmetry of the system [10]. In general, we expect that, for a charge in the system to show anomalous fluctuations, it has to have a vanishing velocity in equilibrium (usually afforded by a $\mathbb{Z}_{2}$ symmetry of the charge) and other modes should have no self-couplings, which would otherwise induce superdiffusion of the charge as in the heat mode with the $5/3$-Levy exponent in anharmonic chains [26]. It is also reasonable to assume that anomalous fluctuations are robust away from charge inertness, but a more precise analysis of such a situation is beyond the scope of the present Letter.
To our knowledge, the above solution is the first exact solution of full BMFT equations in an interacting model and therefore serves as an independent check of their validity. Furthermore, Eqs. (13) and (14) match the solution of the flow equation in ballistic fluctuation theory (see SM), corroborating the identification advocated in Ref. [19].
@@
Conclusions.—We have studied anomalous charge current fluctuations in a family of stochastic cellular automata and characterized both large and typical fluctuations based only on Euler hydrodynamics, reproducing the microscopic results obtained in Ref. [9] when specializing to the single-file (i.e. deterministic) limit $\Gamma=0$. The key insight is that initial fluctuations, carried by Euler hydrodynamics, fully control fluctuations both on the typical and large scales in the single-file case. To implement the idea, we employed BMFT to compute the SCGF and proposed a BMFT-based approach to describe typical fluctuations induced by initial Gaussian fluctuations (in a forthcoming work Ref. [31], it is also shown how BMFT provides the full diffusive-order hydrodynamic equation, where the long-range correlations predicted by BMFT are the source of diffusive-order terms). Our work clarifies for the first time the hydrodynamic origin of anomalous fluctuations, which we believe is generic among systems that exhibit anomalous fluctuations.
While initial fluctuations remain the only source of large fluctuations even in the stochastic case $\Gamma\neq 0$, it turns out that typical fluctuations are nontrivially influenced by the presence of stochasticity, which is not prescribed by initial fluctuations. We expect that stochasticity induces normal diffusion, and it would be very interesting to reproduce Eq. (20) by extending the formalism we introduced to incorporate normal diffusion. The observation of the distribution (19) in a classical spin chain [7] therefore supports the absence of normal diffusion in the easy-axis regime [32, 33]. Other important directions are to obtain the exact FCS of the SCCA via microscopic computations and establish a scattering description of thermodynamics, as recently obtained for classical integrable spin chains [34].
Acknowledgements.—We thank Bruno Bertini, Benjamin Doyon, Sarang Gopalakrishnan, Katja Klobas, and Tomaž Prosen for useful discussions. We are also grateful to Alvise Bastianello for his comments on the manuscript. ŽK is supported by the Simons Foundation as a Junior Fellow of the Simons Society of Fellows (1141511). This paper is dedicated to the memory of Marko Medenjak (1990-2022).

### v2

@@ Abstract
We consider the hydrodynamic origin of anomalous current fluctuations in a family of stochastic charged cellular automata. Using ballistic macroscopic fluctuation theory, we study both typical and large fluctuations of the charge current and reproduce microscopic results which are available for the deterministic single-file limit of the models. Our results indicate that in general initial fluctuations propagated by Euler equations fully characterize anomalous fluctuations on both scales. In the stochastic case, we find an additional contribution to typical fluctuations and conjecture the functional form of the typical probability distribution, which we confirm by numerical simulations.
Introduction.—Statistical physics aims to describe universal behaviors emerging from interactions of a large number of microscopic degrees of freedom. While traditionally the main focus has been equilibrium physics, in recent years it has been appreciated that out-of-equilibrium properties of the system can also display rich universal behaviors. A prominent tool that describes such physics is macroscopic fluctuations theory (MFT) [1], which captures large fluctuations of driven diffusive systems. The validity of MFT has been verified for a variety of systems [2, 3, 4, 5] and also on quantum simulators [6], but large fluctuations in these systems usually behave in a regular way with finite scaled cumulants of observables. This naturally raises the question of whether there exists systems that break such regularity and support new dynamical phenomena.
Recent numerical studies have found dynamical criticality and anomalous spin fluctuations in integrable spin chains [7, 8], which have also been probed experimentally using superconducting qubits [9]. While the mechanism responsible for anomalous spin fluctuations remains to be clarified, an exact calculation of fluctuations in a deterministic charged cellular automaton [10] provided a basic example of dynamical criticality in a deterministic many-body system. The procedure has been generalized to a wider class of single-file dynamics, which was explored in Refs. [11, 12, 13, 14, 15, 16]. Furthermore, anomalous fluctuations have also been observed in the easy-axis regime of an integrable spin chain [8, 17] and other systems including Dirac fluids [18] and stochastic processes [19], hinting at a more general setting supporting anomalous fluctuations.
@@
A system of three-mode hydrodynamic equations similar to Eq. (3) along with the free chiral equations for ϱ±subscriptitalic-$\varrho_{\pm}$ was also recently derived for Dirac fluids by keeping only the most relevant terms based on symmetry considerations in Ref. [18]. In the present case, the hydrodynamic equations are obtained using the exact flux jacobian Eq. (2), and interestingly, are completely insensitive to the crossing probability $\Gamma$. That said, as we will see later, typical fluctuations do turn out to depend nontrivially on $\Gamma$. Before discussing this dependence, we first study large fluctuations of the SCCA, which can be fully characterized using only Euler hydrodynamic data.
@@
As was also observed in Ref. [10], the SCGF Eq. (15) is anomalous in that the scaled higher cumulants $c_{2n>4}=\lim_{t\to\infty}t^{-1}\langle[\hat{J}(t)]^{2n}\rangle^{c}$ are singular. BMFT clarifies the physical origin of this anomalous behavior: since the normal mode $b(x,t)$ of the charge $\varrho_{c}(x,t)$ travels along the ray $x=0$, fluctuations along $x=0$ proliferate and thereby render clustering of connected charge current correlations algebraic in time [27]. While such a scenario generally results in superdiffusion [28], it is avoided here due to the charge conjugation symmetry of the system [11]. In general, we expect that, for a charge in the system to show anomalous fluctuations, it has to have a vanishing velocity in equilibrium (usually afforded by a $\mathbb{Z}_{2}$ symmetry of the charge) and other modes should be independent of their own velocities (i.e. no self-coupling), which would otherwise induce superdiffusion of the charge as in the heat mode with the $5/3$-Levy exponent in anharmonic chains [28]. It is also reasonable to assume that anomalous fluctuations are robust away from charge inertness, but a more precise analysis of such a situation is beyond the scope of the present Letter.
To our knowledge, the above solution is the first exact solution of full BMFT equations in an interacting model and therefore serves as an independent check of their validity. Furthermore, Eqs. (13) and (14) match the solution of the flow equation in ballistic fluctuation theory (see SM) [27, 29], corroborating the identification advocated in Ref. [21].
@@
Conclusions.—We have studied anomalous charge current fluctuations in a family of stochastic cellular automata and characterized both large and typical fluctuations based only on Euler hydrodynamics, reproducing the microscopic results obtained in Ref. [10] when specializing to the single-file (i.e. deterministic) limit $\Gamma=0$. The key insight is that initial fluctuations, carried by Euler hydrodynamics, fully control fluctuations both on the typical and large scales in the single-file case. To implement the idea, we employed BMFT to compute the SCGF and proposed a BMFT-based approach to describe typical fluctuations induced by initial Gaussian fluctuations (in a forthcoming work Ref. [34], it is also shown how BMFT provides the full diffusive-order hydrodynamic equation, where the long-range correlations predicted by BMFT are the source of diffusive-order terms). Following Ref. [18], our work further clarifies the hydrodynamic origin of anomalous fluctuations, which we believe is generic among systems that exhibit anomalous fluctuations.
While initial fluctuations remain the only source of large fluctuations even in the stochastic case $\Gamma\neq 0$, it turns out that typical fluctuations are nontrivially influenced by the presence of stochasticity, which is not prescribed by initial fluctuations. We expect that stochasticity induces normal diffusion, and it would be very interesting to reproduce Eq. (20) by extending the formalism we introduced to incorporate normal diffusion. The observation of the distribution (19) in a classical spin chain [8] therefore supports the absence of normal diffusion in the easy-axis regime [35, 36]. Other important directions would be to obtain the exact FCS of the SCCA via microscopic computations and to establish a scattering description of thermodynamics, as recently obtained for classical integrable spin chains [37].
Acknowledgements.—We thank Bruno Bertini, Benjamin Doyon, Sarang Gopalakrishnan, Enej Ilievski, Katja Klobas, Marcin Mierzejewski, Peter Prelovšek, Tomaž Prosen, and Herbert Spohn for useful discussions. We are also grateful to Alvise Bastianello for his comments on the manuscript. ŽK is supported by the Simons Foundation as a Junior Fellow of the Simons Society of Fellows (1141511). This paper is dedicated to the memory of Marko Medenjak (1990-2022).